Present and future value of a cash flow stream

An investment will pay $100 at the end of each of the next 3 years, $200 at the end of Year 4, $300 at the end of Year 5, and $500 at the end of Year 6.

a. If other investments of equal risk earn 4% annually, what is its present value? Round your answer to the nearest cent.
b.Its future value? Round your answer to the nearest cent

*Please Include how the answer is derived via Financial Calculator, Thanks so much!!
Answer #1

Here's the Present Value Calculation:

Year 1 $100 / (1.0 + .08) = $ 92.59
Year 2 $100 / (1.0 + .08)^2 = $ 85.73
Year 3 $100 / (1.0 + .08)^3 = $ 79.38
Year 4 $200 / (1.0 + .08)^4 = $147.01
Year 5 $300 / (1.0 + .08)^5 = $204.17
Year 6 $500 / (1.0 + .08)^6 = $315.08

Total PV = $923.98

Here's the Future Value calculation:
(Note: since the payment is at the end of each year, the payment for year 1 will compound for only 5 years, the payment for year 2 will compound for 4years, and so on)

Year 1 $100 X (1.0 + .08)^5 = $146.93
Year 2 $100 X (1.0 + .08)^4 = $136.05
Year 3 $100 X (1.0 + .08)^3 = $125.97
Year 4 $200 X (1.0 + .08)^2 = $233.28
Year 5 $300 X (1.0 + .08)^1 = $324.00
Year 6 $500 X (1.0 + .08)^0 = $500.00

Total FV = $1,466.23

answered by: maths